OpenStudy (trisarahtops):

Medal!! Use the graph of f(x) = |x(x^2 − 1)| to find how many numbers in the interval [0.5, 0.75] satisfy the conclusion of the Mean Value Theorem. 1 2 4 None

OpenStudy (trisarahtops):

@Michele_Laino

OpenStudy (nikeboi101):

its answer 1

OpenStudy (trisarahtops):

how do u know?

OpenStudy (nikeboi101):

Did you mean |x(x^2-1)| ? f(x) is continuous for all real x f(x) is differentiable for all x ≠ −1, 0, or 1 f(x) is continuous on interval [1/2, 3/4] f(x) is differentiable on interval (1/2, 3/4) On the interval [1/2, 3/4] f(x) = −x(x^2−1) = −x^3+x f'(x) = −3x^2 + 1 (f(3/4)−f(1/2))/(3/4−1/2) = (21/64−3/8)/(1/4) = −3/16 Now we find all values of x on interval (1/2, 3/4) so that f'(x) = −3/16 −3x^2 + 1 = −3/16 3x^2 = 19/16 x^2 = 19/48 x = √(19/48) = 0.62915287 ---> ok [We ignore negative root, since it is not in the given interval] so yea its answer 1 @trisarahtops

OpenStudy (nikeboi101):

@Trisarahtops

OpenStudy (trisarahtops):

is that right @Michele_Laino

OpenStudy (nikeboi101):

sure double checking is good

OpenStudy (michele_laino):

If I explicit the absolute value, I can write this: \[f\left( x \right) = \begin{array}{*{20}{c}} {x\left( {{x^2} - 1} \right),\quad x \in \left( { - 1,0} \right) \cup \left( {1, + \infty } \right)} \\ {x\left( {1 - {x^2}} \right),\quad x \in \left( { - \infty , - 1} \right) \cup \left( {0,1} \right)} \end{array}\]

OpenStudy (trisarahtops):

so 2?

OpenStudy (michele_laino):

inside the interval \((1/2,3/4)\) we have two definitions for \(f(x)\)

OpenStudy (michele_laino):

so we have to do the computations separately, each computation for each subinterval

OpenStudy (michele_laino):

here is the first computation: \[\frac{{f\left( 1 \right) - f\left( {1/2} \right)}}{{1 - \left( {1/2} \right)}} = 1 - 3{c^2}\] where I have used: \[f\left( x \right) = x\left( {1 - {x^2}} \right)\]

OpenStudy (trisarahtops):

f(1)-f(3/4) / 1-(3/4)

OpenStudy (michele_laino):

after a simple computation I get this: \[c = \sqrt {\frac{{19}}{{48}}} \simeq 0.63 < 1\] as we can see such value of \(c\) is greater than 0.5, and it is less than 1, so it is an acceptable value

OpenStudy (michele_laino):

next we have to do this computation: \[\frac{{f\left( {3/4} \right) - f\left( 1 \right)}}{{\left( {3/4} \right) - 1}} = 1 - 3{c^2}\] where: \[f\left( x \right) = x\left( {1 - {x^2}} \right)\]

OpenStudy (nikeboi101):

so 1 is the answer

OpenStudy (michele_laino):

please wait, I'm doing the remaining computation after a simple computation, I get:

OpenStudy (nikeboi101):

ok

OpenStudy (michele_laino):

I got this: \[c = \sqrt {\frac{{149}}{{128 \cdot 3}}} \simeq 0.623 < 3/4\] being such value less than \(3/4\), it can not be considered an acceptable value, so I confirm the answer of @nikeboi101

OpenStudy (trisarahtops):

Thank you for clearing that up :D

OpenStudy (michele_laino):

:)

OpenStudy (nikeboi101):

c:

OpenStudy (trisarahtops):

you too ;)